# Human as AI Mentor: Enhanced Human-in-the-loop Reinforcement Learning for Safe and Efficient Autonomous Driving

### v3

@@ 5.1.5 Training Environment
We divided the generated driving scenarios into two sets: a training set and a test set. The AV agent was exclusively trained on the training set and subsequently evaluated on the reserved test set. At the outset of each episode, we randomly selected a scenario from either the training set or the test set. Following each training iteration, we deployed the learned agent into the test environment without any human supervision. Experiments are carried out using the MetaDrive/CARLA simulator, and methods are implemented with RLLib (Liang et al., 2018), an efficient distributed learning system.
@@ 5.2.1 Compared to RL and Safe RL Methods
The HAIM-DRL was meticulously evaluated against both traditional RL and Safe RL methods. Traditional RL methods such as SAC (Haarnoja et al., 2018) and PPO (Schulman et al., 2017) were evaluated, with cost applied as an additional negative reward in a process known as reward shaping. In the realm of Safe RL, we scrutinized established benchmarks, including CPO (Achiam et al., 2017), SAC-Lag (Ha et al., 2021), and PPO-Lag (Stooke et al., 2020). The comparative data, as illustrated in Fig. 6 and Table 1, demonstrate that our HAIM-DRL method surpasses these traditional and safety-conscious methods in crucial performance metrics. Notably, due to its underwhelming safety performance, CPO’s results are omitted from Fig. 6 to maintain the focus on more successful methods.
@@ 5.2.2 Compared to IL and Offline RL Methods
We have collected a human demonstration dataset consisting of an hour’s worth of expert demonstrations. This collection translates to about 49K transitions within the training environments  [^3] The high-quality demonstration dataset collected by the human expert in this study is available at: https://github.com/zilin-huang/HAIM-DRL/releases/tag/v0.0.0. This high-quality demonstration dataset boasts a perfect success rate of 100%, with only three instances of disturbance and a notably low safety violation rate of 0.03, thereby establishing a benchmark for our evaluations. Leveraging this dataset, we evaluated passive IL methods such as BC (Bain and Sammut, 1995), active IL methods such as GAIL (Ho and Ermon, 2016), and offline RL methods such as CQL (Kumar et al., 2020).
@@ 5.2.3 Compared to Conventional Human-in-the-loop RL Methods
We benchmarked the performance of two conventional human-in-the-loop methods using the human dataset collected above: HG-DAgger (Kelly et al., 2019) and IWR (Mandlekar et al., 2020). Both of these methods necessitate a preliminary warm-up phase, which involves behavior cloning from a pre-collected dataset. Li et al. (2022c) revealed that lesser quantities, such as 10K or 20K steps of human-collected data, were inadequate to initialize the policy with basic driving skills. Consequently, we employed a more substantial dataset containing 30K transitions for the warm-up phase. Following the initial warm-up phase, both HG-DAgger and IWR integrated human intervention data into their training buffers. Subsequently, they executed behavior cloning to update the policy over 4 epochs. Throughout each epoch, these human-in-the-loop methods necessitated a minimum of 5,000 steps of real-time human supervision.
@@ 5.2.4 Compared to Human-AI Copilot Method
We also conducted a comparative analysis with the recently developed human-AI copilot optimization (HACO) method(Li et al., 2022c). We used the average episodic intervention occurrence to represent the frequency of interventions. As indicated in Fig. 9, at the start of the training, the human expert is more frequently involved in driving demonstrations to prevent the agent from entering hazardous states. Subsequently, as the agent acquires basic driving skills, it tends to select actions more agreeable to the human expert, thereby reducing takeover rate. The takeover cost and takeover rate curves show that both HACO and HAIM-DRL experience a decline in human takeover frequency, while the consistency between agent and human behavior increases. This decline in HAIM-DRL can be attributed to its intervention minimization mechanism, which reduces both the cost and frequency of human takeover.
@@ 5.3.2 Cosine Similarity as Takeover Cost Function
Table 2 (b) displays the outcomes when we altered the takeover cost, as specified in Eq. 23, to a constant value of ‘+1’ in response to a human takeover. It was observed that the agent learned to remain stationary at spawn points, and no movements were recorded during the testing phase. This could be due to the incorrect timing of human takeover, which disrupts the agent’s ability to drive accurately. This is consistent with the findings of  Li et al. (2022c). However, the negative cosine similarity, when used to measure the divergence between the actions of the agent and the human, helps mitigate this issue. This is because the penalty for human takeover is lessened when the agent’s actions align with the human’s takeover.
@@ 5.3.3 The Necessity of Minimizing Explicit Intervention
Table 2 (c) indicates that in the absence of an intervention minimization mechanism, the AV agent is prone to driving straight toward the boundary. This is aligned with the results of  Li et al. (2022c); Wu et al. (2023). This occurs because the agent learns to exploit the human expert’s tendency to take over persistently, thereby inflating the proxy $Q$ values and increasing the cognitive load on the human mentor. Agents might even engage in risky behaviors, such as deliberately deviating from the road when approaching boundaries, to trick the intervention mechanism and compel the human expert to take control and provide demonstrations.
@@ Appendix A Proof of Main Theorem on Explicit Intervention Mechanism
We use the Boolean indicator $I\left(s_{\_}t\right)$ to represent human takeover, and the resulting action applied to the environment is $\hat{a}_{\_}t^{mix}=I\left(s_{\_}t\right)a_{\_}t^{human}+\left(1-I\left(s_{\_}t\right)\right)a_{\_}t^{AV}$ (Wu et al., 2023). As a result, the actual trajectory during the training process is determined by the mixed behavior policy:
@@
Hence, for a given state, we can partition the action space into two categories: actions where human takeover is likely to occur and actions where it is not expected. We refer to the set of actions that are unlikely to be rejected by a human expert at state $s$ as the confident action space, denoted as (Peng et al., 2022):
@@
Therefore, at a given state $s$, the step-wise probability of failure for an arbitrary policy $\pi$ is defined as (Li et al., 2022c):
@@
Moving forward, we introduce the concept of the cumulative discounted probability of failure, denoted as $V_{\_}\pi\left(s_{\_}t\right)$. This is represented mathematically as (Li et al., 2022c):
@@
Lemmas 1 (The performance difference lemma). This lemma was introduced and proven by Kakade and Langford (2002) and serves the purpose of demonstrating the safety of the mixed behavior policy:
@@
Lemmas 2 The cumulative probability of failure of the expert, denoted as $V_{\_}{{\pi_{\_}{human}}}(s)$, is bounded for all state (Li et al., 2022c):

### v5

@@ 5.1.5 Training Environment
We divided the generated driving scenarios into two sets: a training set and a test set. The AV agent was exclusively trained on the training set and subsequently evaluated on the reserved test set. At the outset of each episode, we randomly selected a scenario from either the training set or the test set. Following each training iteration, we deployed the learned agent into the test environment without any human supervision. Experiments are carried out using the MetaDrive/CARLA simulator, and methods are implemented with RLLib [Liang et al., 2018], an efficient distributed learning system.
@@ 5.2.1 Compared to RL and Safe RL Methods
The HAIM-DRL was meticulously evaluated against both traditional RL and Safe RL methods. Traditional RL methods such as SAC [Haarnoja et al., 2018] and PPO [Schulman et al., 2017] were evaluated, with cost applied as an additional negative reward in a process known as reward shaping. In the realm of Safe RL, we scrutinized established benchmarks, including CPO [Achiam et al., 2017], SAC-Lag [Ha et al., 2021], and PPO-Lag [Stooke et al., 2020]. The comparative data, as illustrated in Fig. 6 and Table 1, demonstrate that our HAIM-DRL method surpasses these traditional and safety-conscious methods in crucial performance metrics. Notably, due to its underwhelming safety performance, CPO’s results are omitted from Fig. 6 to maintain the focus on more successful methods.
@@ 5.2.2 Compared to IL and Offline RL Methods
We have collected a human demonstration dataset consisting of an hour’s worth of expert demonstrations. This collection translates to about 49K transitions within the training environments  [^3] The high-quality demonstration dataset collected by the human expert in this study is available at: https://github.com/zilin-huang/HAIM-DRL/releases/tag/v0.0.0. This high-quality demonstration dataset boasts a perfect success rate of 100%, with only three instances of disturbance and a notably low safety violation rate of 0.03, thereby establishing a benchmark for our evaluations. Leveraging this dataset, we evaluated passive IL methods such as BC [Bain and Sammut, 1995], active IL methods such as GAIL [Ho and Ermon, 2016], and offline RL methods such as CQL [Kumar et al., 2020].
@@ 5.2.3 Compared to Conventional Human-in-the-loop RL Methods
We benchmarked the performance of two conventional human-in-the-loop methods using the human dataset collected above: HG-DAgger [Kelly et al., 2019] and IWR [Mandlekar et al., 2020]. Both of these methods necessitate a preliminary warm-up phase, which involves behavior cloning from a pre-collected dataset. Li et al. [2022c] revealed that lesser quantities, such as 10K or 20K steps of human-collected data, were inadequate to initialize the policy with basic driving skills. Consequently, we employed a more substantial dataset containing 30K transitions for the warm-up phase. Following the initial warm-up phase, both HG-DAgger and IWR integrated human intervention data into their training buffers. Subsequently, they executed behavior cloning to update the policy over 4 epochs. Throughout each epoch, these human-in-the-loop methods necessitated a minimum of 5,000 steps of real-time human supervision.
@@ 5.2.4 Compared to Human-AI Copilot Method
We also conducted a comparative analysis with the recently developed human-AI copilot optimization (HACO) method[Li et al., 2022c]. We used the average episodic intervention occurrence to represent the frequency of interventions. As indicated in Fig. 9, at the start of the training, the human expert is more frequently involved in driving demonstrations to prevent the agent from entering hazardous states. Subsequently, as the agent acquires basic driving skills, it tends to select actions more agreeable to the human expert, thereby reducing takeover rate. The takeover cost and takeover rate curves show that both HACO and HAIM-DRL experience a decline in human takeover frequency, while the consistency between agent and human behavior increases. This decline in HAIM-DRL can be attributed to its intervention minimization mechanism, which reduces both the cost and frequency of human takeover.
@@ 5.3.2 Cosine Similarity as Takeover Cost Function
Table 2 (b) displays the outcomes when we altered the takeover cost, as specified in Eq. 23, to a constant value of ‘+1’ in response to a human takeover. It was observed that the agent learned to remain stationary at spawn points, and no movements were recorded during the testing phase. This could be due to the incorrect timing of human takeover, which disrupts the agent’s ability to drive accurately. This is consistent with the findings of  Li et al. [2022c]. However, the negative cosine similarity, when used to measure the divergence between the actions of the agent and the human, helps mitigate this issue. This is because the penalty for human takeover is lessened when the agent’s actions align with the human’s takeover.
@@ 5.3.3 The Necessity of Minimizing Explicit Intervention
Table 2 (c) indicates that in the absence of an intervention minimization mechanism, the AV agent is prone to driving straight toward the boundary. This is aligned with the results of  Li et al. [2022c], Wu et al. [2023]. This occurs because the agent learns to exploit the human expert’s tendency to take over persistently, thereby inflating the proxy $Q$ values and increasing the cognitive load on the human mentor. Agents might even engage in risky behaviors, such as deliberately deviating from the road when approaching boundaries, to trick the intervention mechanism and compel the human expert to take control and provide demonstrations.
@@ Appendix A Proof of Main Theorem on Explicit Intervention Mechanism
We use the Boolean indicator $I\left(s_{\_}t\right)$ to represent human takeover, and the resulting action applied to the environment is $\hat{a}_{\_}t^{mix}=I\left(s_{\_}t\right)a_{\_}t^{human}+\left(1-I\left(s_{\_}t\right)\right)a_{\_}t^{AV}$ [Wu et al., 2023]. As a result, the actual trajectory during the training process is determined by the mixed behavior policy:
@@
Hence, for a given state, we can partition the action space into two categories: actions where human takeover is likely to occur and actions where it is not expected. We refer to the set of actions that are unlikely to be rejected by a human expert at state $s$ as the confident action space, denoted as [Peng et al., 2022]:
@@
Therefore, at a given state $s$, the step-wise probability of failure for an arbitrary policy $\pi$ is defined as [Li et al., 2022c]:
@@
Moving forward, we introduce the concept of the cumulative discounted probability of failure, denoted as $V_{\_}\pi\left(s_{\_}t\right)$. This is represented mathematically as [Li et al., 2022c]:
@@
Lemmas 1 (The performance difference lemma). This lemma was introduced and proven by Kakade and Langford [2002] and serves the purpose of demonstrating the safety of the mixed behavior policy:
@@
Lemmas 2 The cumulative probability of failure of the expert, denoted as $V_{\_}{{\pi_{\_}{human}}}(s)$, is bounded for all state [Li et al., 2022c]:
@@ Appendix B Hyper-parameters
Zilin Huang received his B.S. degree from School of Electromechanical Engineering, Guangdong University of Technology in 2018. He received his M.S. degree in Communication and Transportation Engineering from South China University of Technology in 2021. He is currently pursuing a Ph.D. degree at the Department of Civil and Environmental Engineering, University of Wisconsin-Madison, USA. Before joining UW-Madison, he worked at the Center for Connected and Automated Transportation (CCAT), Purdue University, USA. His research interests include human-centered AI, autonomous driving, robotics, human-AI interaction, intelligent transportation.
Zihao Sheng received his B.S. degree in Automation from Xi’an Jiaotong University, Xi’an, China in 2019 and M.S. degree in Control Engineering from Shanghai Jiao Tong University, Shanghai, China in 2022. He is currently pursuing a Ph.D. degree at the Department of Civil and Environmental Engineering, University of Wisconsin-Madison, USA. His main research interests include human-centered AI, autonomous driving, reinforcement learning, and intelligent transportation.
Chengyuan Ma received the B.S. and Ph.D. degrees in traffic engineering from Tongji University in 2017 and 2023, respectively. He is currently a Postdoctoral Research Associate with the Department of Civil and Environmental Engineering of the University of Wisconsin-Madison. His research interests include traffic control with connected autonomous vehicles and transportation big data analytics.
Sikai (Sky) Chen is an Assistant Professor at the Department of Civil and Environmental Engineering, University of Wisconsin-Madison. He received his Ph.D. in Civil Engineering with a focus on Computational Science & Engineering from Purdue University. His research centers around three major themes: human users, AI, and transportation. He aims to innovate and develop safe, efficient, sustainable, and human-centered transportation systems using cutting-edge methods and technologies. The focus is on incorporating human behaviors, interactive autonomy, and intelligent control frameworks. In addition, he is a member of two ASCE national committees: Connected & Autonomous Vehicle Impacts, and Economics & Finance; and TRB Standing Committee on Statistical Methods (AED60). More information can be found at Sky-Lab: https://sky-lab-uw.github.io/